A box contains two black balls and three gold balls. Two
balls are randomly drawn in succession from the box.
a. If there is no replacement, what is the probability
that both balls are black?
b. If there is replacement before the second draw, what
is the probability that both balls are black?
I am having a hard time putting these into formulas. The answers I got are a. 2/3 and b. 2/5

Found 2 solutions by gerrylucas, stanbon:
Answer by gerrylucas(2)

I believe the answer is a 1/20 and b 1/25

Answer by stanbon(75887)

A box contains two black balls and three gold balls. Two
balls are randomly drawn in succession from the box.
a. If there is no replacement, what is the probability
that both balls are black?
Ans: P(black and black) = P(black)*P(black | black)
= (2/5)(1/4) = 2/20 = 1/10
-------------------------------------
b. If there is replacement before the second draw, what
is the probability that both balls are black?
Ans: P(black and black) = (2/5)(2/5) = 4/25
------------------------------------------
Cheers,
Stan H.